A uniform distribution is defined over the interval from 10 to 22. Find the mean and the standard deviation of uniform distribution. Find the probability of value between 18 and 20.
Answer #1

The mean of the uniform distribution= (10+22) / 2

= 32/2

= 16

The standard deviation of the uniform distribution= (22-10)2 / 12

= 144/12

=12

The probability that the value is between 18 and 20= (20-18) / (22-10)

= 2/12

= 1/6

= 0.167 ~ 16.7%